THE GOVERNMENT HAS $100 MILLION TO SPEND ON THE ECONOMY, BUT IS UNCERTAIN IF TO DO: OUTRIGHT GOVERNMENT EXPENDITURES OR A TAS CUT; ASSUMING AN MPC = 0.75 AND A TAX RATE = 0.05, WHICH OPTION SHOULD THE GOVERNMENT SELECT THAT WILL MAXIMIZE IMPACT ON THE ECONOMY?
1
Expert's answer
2020-12-10T09:44:22-0500

Government spending multiplier is greater than tax multiplier so in order to have a greater impact government should raise its spending.


Spending multiplier "= 1/(1 - MPC(1-t)) = 1/(1-0.75 \times 0.95) = 3.478"

Tax multiplier "= -MPC(1-t)\/(1 - MPC(1-t)) = -0.7125\/0.2875 = -2.478"

Hence rise in GDP with increase in "G = 100 \times 3.478 = 347.8" billion

Rise in GDP with tax cut "= -100 \times -2.478 = 247.8" billion